The ratio of the measures of three sides of a triangle are 9:8:7, and its perimeter is 144. What is the measure of each of the 3 sides of the triangle?
Answer by checkley77(12844)

9n+8n+7n=144
24n=144
n=144/24
n=6 ans. for the multiplier.
9*6=54 for side 1
8*6=48 for side 2
7*6=42 for side 3
Proof:
54+48+42=144
144=144